OpenStudy (anonymous):

how do i find a formula for a function " f " that satisfies certain (infinite) limit conditions?

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0}=(f(x+\Delta x)−f(x))/\Delta x\]

OpenStudy (anonymous):

this is an example of what i am asking about

OpenStudy (anonymous):

#49

OpenStudy (zarkon):

look for a rational function that works. It is not a complicated function end the end

OpenStudy (kmeis002):

Recall That \[\lim_{x \rightarrow \pm \infty} \frac{ax^{n}+...}{bx^{m}+...} \] Where they are rational functions, recall that so long as m>n, then the limit at infinities will equal zero. The other conditions depend on the denominator and numerator, respectively. So try to come up with a vertical asymptote at x = 0 (what does that mean for the denominator?) And how would you get f(2) = 0? (think about the numerator)

OpenStudy (anonymous):

thank you soo much :)))

OpenStudy (anonymous):

i found the answer but i have one question :/

OpenStudy (anonymous):

the purple box is the real answer, but i dont know where the negative (-) comes from

OpenStudy (kmeis002):

Do you mean the 2-x?

OpenStudy (anonymous):

yeah

OpenStudy (kmeis002):

For the condition \[ f(2) = 0\] Implies the numerator must be zero so it must have a factor that = 0. When x = 2 then (2-x) = 0

OpenStudy (anonymous):

but doesnt (x-2) also equal two

OpenStudy (anonymous):

they both equal to 2

OpenStudy (kmeis002):

Ah, it arises from the fact that \[ \lim_{x \rightarrow 0} f(x) = -\infty\] (x-2) when x approaches zero results in a negative (2-x) when x approaches zero results in a positive And since (x-3) results in a negative, your factor up top must result in a positive or else \[ \lim_{x \rightarrow 0} f(x) = \infty\]

OpenStudy (anonymous):

OH OKAY! thank you soo much